Champion Bakers uses specialized ovens to bake its bread. One oven costs $840,000 and lasts about 3 years before it needs to be replaced. The annual operating cost per oven is $10,000. What is the equivalent annual cost of an oven if the required rate of return is 11 percent? (Round your answer to whole dollars)

$-359,381

$-348,281

$-864,437

$-353,739
Answer #1

Present value of operating costs = 10,000 * PVAF( 11% , 3 years) = 10,000*2.44371 = 24437.1

Total of the present value = 840,000 + 24437.1 = 864437.1

EAC = 864437.1 / PVAF of $1( 11%, 3 years) = 864437.1 /2.44371 = 353,739

Option $- 353,739